Operations Report – May 5th

Crew 193 Operations Report 05May2018

SOL over

Name of person filing report: David Attig

Non-nominal systems: Two piece suits still need straps adjusted for neck ring. Generator/inverter/charger. One piece suit #9
Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now).

Generator (hours run): 8:45 pm to 7:30 am.

Solar— SOC 5 % (Before generator is run last night)

Diesel – 60 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0 gallons.

Water (trailer) – 500 gallons.

Water (static) – 500 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134635.2

Toilet tank emptied: no
Date last emptied: May 2

ATV’s (Honda, 300, 350.1, 350.2, 350.3): No

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: 350.2 is out of service, due to a previous issue. Honda has a hole in the seat. 300 would not start, there is a grinding or clicking sound (dead battery?)

Deimos rover used: No
Hours: 125.6
Beginning charge: 100%
Ending charge: N/A
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes
Hours: 36.0
Beginning charge: 100 %
Ending charge: 99 %
Currently charging: Yes

Opportunity rover used: Yes
Hours: 31.6
Beginning charge: 100 %
Ending charge: 98 %
Currently charging: Yes

Curiosity rover used: Yes
Hours: 29.9
Beginning charge: 100 %
Ending charge: 98 %
Currently charging: No

Hab Car used?: Yes
Where was Hab Car used?: gas station
Why was Hab Car used?: Water replenishment

General notes and comments: none

Summary of internet: We’re out, same as normal.
Summary of suits and radios: Two piece suits still need strap maintenance

Summary of Hab operations: Super nominal.

Summary of GreenHab operations: Water saturated

Summary of ScienceDome operations: Concluded

Summary of RAM operations: cleaned

Summary of health and safety issues: None to report

Questions, concerns and requests to Mission Support: Sol over

Copyright © The Mars Society. All rights reserved. | Main Site